# Shaping chaos in bilayer graphene cavities

### v2

@@ III.1 Level Statistics
The rotated cavity shows a qualitatively different behavior. As illustrated by Fig. 2(c,d) for $\theta=15^{\circ}$, the AA sector remains intermediate with $\langle\tilde{r}\rangle=0.516$, while the $E$ sector shifts toward Wigner–Dyson correlations with $\langle\tilde{r}\rangle=0.558$ and a strongly suppressed exponential tail. The data are well described by a mixed GOE–GUE interpolation [mehta2004random](see Materials and Methods for the fitting form), demonstrating that lattice–boundary misalignment removes the near-integrable channel present at $\theta=0^{\circ}$ and drives the spectrum toward fully chaotic statistics. The rigidity shows a consistent trend: $\langle\Delta_{3}\rangle$ decreases to $0.31$ in AA and $0.24$ in $E$, approaching the GOE/GUE limits.
Figure 3: Angle dependence of $\langle\tilde{r}\rangle$ and $\langle\Delta_{3}\rangle$ on $\theta$. (a) and (b) present the dependence of the $r$-value and the average spectral rigidity on the rotation angle $\theta$ for the AA and $E$ subspaces, respectively. In the AA subspaces, two lines are shown: for the unrotated cases $\theta=0^{\circ}$ and $30^{\circ}$, the blue solid line corresponds to the pseudointegrable sector, while the purple dashed line represents the integrable sector.
Fig. 3 summarizes the evolution with rotation angle. In the AA sector, one branch remains in the intermediate (semi-Poisson-like) regime across all angles, while the branch that is Poisson-like at $\theta=0^{\circ}$ and $30^{\circ}$ rapidly increases toward Wigner–Dyson values as $\theta$ departs from these commensurate alignments. In the $E$ sector, $\langle\tilde{r}\rangle$ increases more steadily and approaches the GOE/GUE range over a broad interval of angles. The averaged rigidity $\langle\Delta_{3}\rangle$ exhibits the same overall trend, shown in Fig. 3(b), indicating that the enhancement of chaotic correlations is robust and does not require fine tuning of $\theta$.
@@ III.2 Eigenstates
The change in the correlation length statistics upon rotation is striking. We consider a cavity rotated by $\theta=15^{\circ}$, as shown in Figs. 4(c4) and (c5). For both the AA and $E$ sectors, the average correlation length decreases dramatically compared to the unrotated case, reaching $\langle l\rangle_{A}=39.2\,\mathrm{nm}$ and $\langle l\rangle_{E}=23.3\,\mathrm{nm}$. Notably, in the $E$ subspace, the correlation lengths of most eigenstates cluster around the wavelength $\lambda=20\,\mathrm{nm}$, indicating the emergence of spatially uncorrelated, random-wave–like eigenstates.

### v3

@@ III.1 Level Statistics
The rotated cavity shows a qualitatively different behavior. As illustrated by Fig. 2(c,d) for $\theta=15^{\circ}$, the $A$ sector remains intermediate with $\langle\tilde{r}\rangle=0.516$, while the $E$ sector shifts toward Wigner–Dyson correlations with $\langle\tilde{r}\rangle=0.558$ and a strongly suppressed exponential tail. The data are well described by a mixed GOE–GUE interpolation [mehta2004random](see Materials and Methods for the fitting form), demonstrating that lattice–boundary misalignment removes the near-integrable channel present at $\theta=0^{\circ}$ and drives the spectrum toward fully chaotic statistics. The rigidity shows a consistent trend: $\langle\Delta_{3}\rangle$ decreases to $0.31$ in $A$ and $0.24$ in $E$, approaching the GOE/GUE limits.
Figure 3: Angle dependence of $\langle\tilde{r}\rangle$ and $\langle\Delta_{3}\rangle$ on $\theta$. (a) and (b) present the dependence of the $r$-value and the average spectral rigidity on the rotation angle $\theta$ for the $A$ and $E$ subspaces, respectively. In the AA subspaces, two lines are shown: for the unrotated cases $\theta=0^{\circ}$ and $30^{\circ}$, the blue solid line corresponds to the pseudointegrable sector, while the purple dashed line represents the integrable sector.
@@
Fig. 3 summarizes the evolution with rotation angle. In the $A$ sector, one branch remains in the intermediate (semi-Poisson-like) regime across all angles, while the branch that is Poisson-like at $\theta=0^{\circ}$ and $30^{\circ}$ rapidly increases toward Wigner–Dyson values as $\theta$ departs from these commensurate alignments. In the $E$ sector, $\langle\tilde{r}\rangle$ increases more steadily and approaches the GOE/GUE range over a broad interval of angles. The averaged rigidity $\langle\Delta_{3}\rangle$ exhibits the same overall trend, shown in Fig. 3(b), indicating that the enhancement of chaotic correlations is robust and does not require fine tuning of $\theta$.
@@ III.2 Eigenstates
The change in the correlation length statistics upon rotation is striking. We consider a cavity rotated by $\theta=15^{\circ}$, as shown in Figs. 4(c4) and (c5). For both the $A$ and $E$ sectors, the average correlation length decreases dramatically compared to the unrotated case, reaching $\langle l\rangle_{A}=39.2\,\mathrm{nm}$ and $\langle l\rangle_{E}=23.3\,\mathrm{nm}$. Notably, in the $E$ subspace, the correlation lengths of most eigenstates cluster around the wavelength $\lambda=20\,\mathrm{nm}$, indicating the emergence of spatially uncorrelated, random-wave–like eigenstates.